OpenStudy (anonymous):

Can the following equation be solved by separation of the variables? y''+3y'+y = 0

OpenStudy (michele_laino):

Sorry, I don't think

OpenStudy (michele_laino):

Please try this substitution: \[y(x)=k*e ^{\lambda x}\] where k is a constant, and \[\lambda \] is another constant to be determined

OpenStudy (michele_laino):

Please try the substitution tha I have wrote and write the equation which you got after that substitution, note that the variable, in your re-writed equation, must be \[\lambda \]

OpenStudy (anonymous):

I get\[y(x) = a \exp(\frac{ -1 }{ 2 }(\sqrt{5}+3)x) +b \exp(\frac{ 1 }{ 2 }(\sqrt{5}-3)x )\]

OpenStudy (michele_laino):

that's right!